OpenStudy (anonymous):

Solve for x. x minus three over two equals negative five A.–16 B.–4 C.–13 D.–7

OpenStudy (anonymous):

This is what I mean

OpenStudy (anonymous):

x-3/2=-5 First multiple both sides of the equation by 2 in order to eliminate the denominator of x-3/2. x-3/2*2=-5*2 x-3=-10 Now please try to solve the equation yourself.

OpenStudy (anonymous):

NiallsMofo1019?

OpenStudy (anonymous):

Thx @Chineseboy15 :)